OpenStudy (anonymous):

5r2 −44r+120=−30+11r

OpenStudy (anonymous):

\[5r ^{2}-44r+120=-30+11r\] \[5r ^{2} -44r - 11r +120 +30 = 0\] \[5r ^{2} - 55r + 150 = 0\] divide by 5 \[r ^{2} - 11r + 30 = 0\] Now Factorize

OpenStudy (anonymous):

\[r ^{2} - 5r-6r+30=0\] \[r(r-5) -6(r-5) = 0\] \[(r-5)(r-6) = 0\]